7. Scores on a recent national Mathematics exam were normally distributed with a mean of 82 and a standard deviation of 7.

A. What is the probability that a randomly selected exam score is less than 70

B. What is the probability that a randomly selected exam score is greater than 90?

C. If the top 2.5% of test scores receive Merit awards, what is the lowest score necessary to receive a merit award?
Answer #1

Solution :

Given that ,

A.

P(x < 70) = P[(x - ) / < (70 - 82 ) / 7]

= P(z < -1.71)

= 0.0436

Probability = 0.0436

B.

P(x > 90) = 1 - P(x < 90)

= 1 - P[(x - ) / < (90 - 82) / 7)

= 1 - P(z < 1.14)

= 1 - 0.8729

= 0.1271

Probability = 0.1271

C.

Using standard normal table ,

P(Z > z) = 2.5%

1 - P(Z < z) = 0.025

P(Z < z) = 1 - 0.025

P(Z < 1.96) = 0.975

z = 1.96

Using z-score formula,

x = z * +

x = 1.96 * 7 + 82 = 95.72

lowest score = 95.72